OpenStudy (anonymous):

how do i calculate the heat of dissolution of a salt in cal/g and in Kcal/mol? salt NH(4)NO(3): mass of water=97.52g mass of salt=10.0g Ti of water and salt=22 degrees celsius Tf of solution=14 degrees celsius

OpenStudy (callisto):

Heat absorbed = mcΔT = 97.52 (4.2) (22-14) = 3276.762 J what is cal?? I'm not sure if the following is correct heat of dissolution of a salt in cal/g = 3276.762 J / 10.0 g = work it out yourself heat of dissolution of a salt in Kcal/mol = 3276.762 J/ (10/80) = 3276.762 / 0.125 = work it out yourself

OpenStudy (jfraser):

1 calorie is 4.18J, so divide your J by 4.18 to get cal.

OpenStudy (callisto):

Oh, thanks for the information! i've never do question involved cal in chemistry before, so i don't know it...